Bert is twice victor's age now. After 5 years he will be 1 1/2 times older than victor. How old is each of them now?
Answer by macston(5194)

.
B=Bert's age; V=Victor's age
.
B=2V . Use this to substitute for B below.
.
B+5=1.5(V+5) . Substitute for B from above.
2V+5=1.5V+7.5
0.5V=2.5
V=5
ANSWER 1: Victor is 5 years old.
B=2V=2(5)=10
ANSWER 2: Bert is 10 years old.
.
CHECK:
B+5=1.5(V+5)
10+5=1.5(5+5)
15=1.5(10)
15=15